A truck and a car leave an intersection at the same time traveling in the same direction. The
truck’s speed is 30 mph, and the speed of the car is 35 mph. After how long will the car be 5 miles ahead of the truck?

Answer by oberobic(2304)

d = r*t, where d=distance, r=rate, and t=time.
so
t = d/r
.
Truck:
r = 30 mph
d = x miles
t = x/30
.
Car:
r = 35 mph
d = x+5 miles
t = (x+5)/35
.
t = t
so
x/30 = (x+5)/35
.
Cross multiply.
.
35x = 30(x+5)
35x = 30x + 150
5x = 150
x = 30 miles
.
t = 30 miles/30 mph = 1 hr
@ 30 mph it takes 1 hr to travel 30 miles
.
Checking
t = 35 miles/35 mph = 1 hr
@ 35 mph it takes 1 hr to travel 35 miles
.
So, the car will be 5 miles ahead in 1 hr.
.
Done.